OpenStudy (calculusxy):

MEDAL!!! The sum of the digits of a certain two-digit number is 8. When you reverse its digits you decrease the number by 18. Find the number. (Solve this using elimination.)

OpenStudy (calculusxy):

@SolomonZelman

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \underline{b}\underline{a} =\underline{a}\underline{b} +18 }\) \(\color{#000000 }{ \displaystyle a+b=8 }\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \underline{b}\underline{a} =\underline{a}\underline{b} +18 }\) can be also written as \(\color{#000000 }{ \displaystyle 10b+a=10a+b+18 }\) And with the \(a+b=8\) you get, \(\color{#000000 }{ \displaystyle 10b+a=10a+b+18 }\) \(\color{#000000 }{ \displaystyle a+b=8 }\)

OpenStudy (solomonzelman):

So you just have a system of equations to solve (hope it's clear that when I underline I mean digits)

OpenStudy (solomonzelman):

I came up with the correct answer for the two digit number and the inversed-digit number, but I have to give you a chance

OpenStudy (anonymous):

Solve \[x+y=8\] and \[10x + y - 18 = x + 10y\]\[9x-9y = 18\]\[x-y=2\]Add the two equations gives\[2x=10\]\[x=5\]Therefore, y=3, and the number is 53.

OpenStudy (solomonzelman):

you are not supposed to give direct answers, this is against the site's policy...

OpenStudy (anonymous):

I'm well aware. My apologies, I assumed this was a challenge.

OpenStudy (solomonzelman):

Well, I didn't see the question say "challenge", but I suppose this time its fine; I'm not a moderator anyway:)

OpenStudy (solomonzelman):

Also the number is not 53 (look at the question)

OpenStudy (calculusxy):

@SolomonZelman I still don't understand how you got your equations.

OpenStudy (anonymous):

I warn others about providing direct answers ALL the time.

OpenStudy (solomonzelman):

Ok, do you agree that 38 can be recorded as 30•10+8?

OpenStudy (calculusxy):

Yes

OpenStudy (solomonzelman):

Ok, and 83 can be recorded as 8•10+3, right?

OpenStudy (anonymous):

I beg to differ on the correctness of my answer. The sume of the digits of 53 is eight and when you reverse its digits, i.e. 35, you decrease the number by 18.

OpenStudy (solomonzelman):

you are looking for the original number, not when the digits are revered. So 35, not 53.

OpenStudy (anonymous):

Try again.

OpenStudy (solomonzelman):

calculusxy, don't want to bug.... u there?

OpenStudy (calculusxy):

Yes I agree that 83 can be recorded as 8•10+3.

OpenStudy (solomonzelman):

Ok, now back to the problem, The sum of the digits of a certain two-digit number is 8. When you reverse its digits you decrease the number by 18. Find the number. (Solve this using elimination.) So let's call our digits a and b.

OpenStudy (calculusxy):

ok

OpenStudy (solomonzelman):

Our original number is \(\underline{a}\underline{b}\) (I am underlining to denote that they are digits, not a•b)

OpenStudy (solomonzelman):

\(\underline{a}\underline{b}=a\cdot 10+b\) alternatively, like that we can record our original number.

OpenStudy (solomonzelman):

We are given that when the digits are reversed, the number increases by 18, or in equation world that is, \(\underline{b}\underline{a}=18+\underline{a}\underline{b}\) (You need to add 18 to original number to make it = to the number with reversed digits -- that is the number with revered digits is +18 greater)

OpenStudy (solomonzelman):

And we also know that the sum of digits is 8, \(a+b=8\)

OpenStudy (calculusxy):

the number is decreased by 18

OpenStudy (anonymous):

Question states the number is decreased by 18 when the digits are reversed, not increased.

OpenStudy (solomonzelman):

I apologize @ospreytriple :)

OpenStudy (solomonzelman):

So when the original number is decreased by 18, we have \(\underline{b}\underline{a}=\underline{a}\underline{b}-18\)

OpenStudy (anonymous):

You've got it well under control. Sorry for my earlier indiscretion.

OpenStudy (solomonzelman):

calculusxy, are you lost?

OpenStudy (calculusxy):

no you can proceed

OpenStudy (calculusxy):

but i am confused as to how can i solve it when there is not a certain result to the equation (numerically)?

OpenStudy (solomonzelman):

\(\underline{a}\underline{b}\) - original number \(\underline{b}\underline{a}\) - number with reversed digits sum of digits is 8, - \(a+b=8\) Alternatively, \(10a+b\) - original number \(10b+a \) - number with reversed digits When you reverse digits number decreases by 18, \(\underline{a}\underline{b}-18=\underline{b}\underline{a}\) or alternatively, \(10a+b-18=10b+a \)

OpenStudy (solomonzelman):

Therefore, you have your system of equations; \(10a+b-18=10b+a \) \(a+b=8 \)

OpenStudy (solomonzelman):

questions about this so far?

OpenStudy (calculusxy):

wait. i am just looking over what you just posted.

OpenStudy (solomonzelman):

:) take your time

OpenStudy (calculusxy):

where did you get 10b + a?

OpenStudy (calculusxy):

or how did you get 10b + a ?

OpenStudy (solomonzelman):

82 can be recorded as 8•10+2, right?

OpenStudy (calculusxy):

yes

OpenStudy (solomonzelman):

and 28 can also be recorded as 2•10+8. Correct?

OpenStudy (calculusxy):

oh okay

OpenStudy (solomonzelman):

This is how I am getting 10a+b 10b+a

OpenStudy (solomonzelman):

10a+b, is the original number \(\underline{a}\underline{b}\) 10b+a, is the reversed-digits number \(\underline{b}\underline{a}\)

OpenStudy (calculusxy):

i tried to simplify the first equation \(10a + b - 18 = 10b + a\) into \(-18 = 9b - 9a\).

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 10a+b-18=10b+a }\) \(\color{#000000 }{ \displaystyle -18=9b-9a }\) And you know that sum of digits is 8 -- i.e. \(a+b=8\)

OpenStudy (calculusxy):

so i can use the system of equations: \(-18 = 9b - 9a\) \(a + b = 8\)

OpenStudy (solomonzelman):

yes, you can solve this system of equations

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle -18=9b-9a }\) \(\color{#000000 }{ \displaystyle 8=b+a }\) \(\color{#000000 }{ \displaystyle -18=9b-9a }\) \(\color{#000000 }{ \displaystyle 72=9b+9a }\) and then add the equations (one way to do this)

OpenStudy (calculusxy):

a = 5 b = 3 number is 53?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 54=18b }\) \(\color{#000000 }{ \displaystyle b=3}\) a+b=8 ---> a+3=8 ----> a=5

OpenStudy (solomonzelman):

Yes, you are correct

OpenStudy (calculusxy):

thank you!

OpenStudy (solomonzelman):

So the original number (the one requested) is 53.

OpenStudy (solomonzelman):

yw